# Optimal Control on Positive Cones

### v1

@@ 3.2 A Linear Regulator on a Polyhedral Cone
In this subsection we consider the application of 1 when $\mathcal{P}$ is the polyhedral cone (which are always positive and proper)
@@
When minimising the cost in 1 ($i$), we are in effect optimizing performance under the constraint that the control keeps the system state positive. In what follows will recover the results of [5, Theorem 1].
@@
It is readily checked that 1) amounts to requiring that $A\geq{}\left|{EB}\right|$. For 2) and 3) it is convenient to appeal to 2. Condition ($ii$) of this statement postulates that the set
$\mathcal{C}_{\mu}=\left\{\lambda:\left(\lambda,\mu\right)\in\mathcal{P}^{*}\right\}$
has a minimum element. From eq. 9 we see that for any $\mu$, $y$ and $z$ can be chosen according to
@@
where $w\geq{}0$. The minimum element is therefore given by $\lambda^{*}=E^{\top}\left|{\mu}\right|$. This implies that for (2) we require that $s\geq{}E^{\top}\left|{r}\right|$, and also that
@@
3.3 Linear Quadratic Regulation with Additional Structure
@@
We now have the following theorem, concerning the matrix LQR problem from section 3.1. The first part of the statement sets up the LQR problem for systems that preserve both the semi-definite cone, and $\mathcal{P}_{Q}$. The important part of the statement is eq. 10. In words, this equation shows that whenever the LQR problem is well defined, the optimal control can be computed through a state-feedback written in terms of matrices structured by the dual cone $\mathcal{P}_{Q}$, revealing the extra structure in the control law.
Theorem 3.
Let $\mathcal{P}_{Q}$ be as in 1, and suppose that $F$, $G$, $S$, $R_{1}$ and $R_{2}$ are such that:
  1. (a)
    $\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\in\mathcal{P}_{Q}$ implies that for some $V_{1}$ and $V_{2}$,
    $\begin{bmatrix}FXF^{\top}+FU_{1}G^{\top}+GU_{1}^{\top}F^{\top}+GU_{2}G^{\top}&V_{1}\\ V_{1}^{\top}&V_{2}\end{bmatrix}\in\mathcal{P}_{Q}.$
  2. (b)
    $\begin{bmatrix}S&R_{1}\\ R_{1}^{\top}&R_{2}\end{bmatrix}\in\mathcal{P}_{Q}^{*}.$
Then whenever the optimization
over $\{U_{1}(t),U_{2}(t)\}_{t=0}^{\infty},$
$\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\succeq{}0,X(0)\succeq{}0$
has a finite value, there exists a
$\begin{bmatrix}\Lambda{}&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\in\mathcal{P}_{Q}^{*}$
such that the input sequence $\{\overline{U}_{1}(t),\overline{U}_{2}(t)\}_{t=0}^{\infty}$ that minimizes the cost satisfies
$\begin{bmatrix}\Lambda{}&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\begin{bmatrix}X(t)&\overline{U}_{1}(t)\\ \overline{U}_{1}(t)^{\top}&\overline{U}_{2}(t)\end{bmatrix}=0$ (10)
for all $t$.
@@
To illustrate the theorem, suppose that the matrices $F$ and $G$ have a block-wise circulant structure. That is,
@@
It follows from these expressions that for such matrices, 3(a) holds. As is shown in 1 in the Appendix A,
$\mathcal{P}_{Q}^{*}=\left\{T^{\top}A\left(Y\right)T:Y\succeq{}0\right\},$
where $A$ and $T$ are as in 1. That is, the dual cone consists of the positive semi-definite matrices with blocks that have a similar circulant structure to the matrices $F$ and $G$. 3 then states that provided the optimization criteria is chosen appropriately, the matrix that encodes the optimal control law eq. 10 will have the same block-wise circulant structure. Interestingly this cone based treatment makes no use of any orthogonality properties of the matrix $Q$ allowing similar insights to be derived for general dynamics that take the form of eq. 11, see for example [10].
@@ 4 Conclusions
We have proved a general result on optimal control on positive cones, covering standard linear quadratic control and corresponding results for positive systems as special cases. Of course, many other special cases can be derived by defining other relevant cones, or combinations of the ones above. Hopefully, this will in the future form the basis for a powerful toolbox combining the versatility of classical linear quadratic control with the rich scalability properties of positive systems and shortest path problems in networks. Preliminary steps in this direction have already been taken in [11] and [12].
Appendix A Proof of 3
The result will follow from an application of 1, based on the following simple properties of $\mathcal{P}_{Q}$.
Lemma 1.
Let $\mathcal{P}_{Q}$ be as in 1. Then:
  1. ($i$)
    If $P\succeq{}0$, then $P\in\mathcal{P}_{Q}$.
  2. ($ii$)
    $\begin{bmatrix}\Lambda&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\in\mathcal{P}^{*}$
    only if
    $\begin{bmatrix}M_{1}M_{2}^{+}M_{1}^{\top}&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\in\mathcal{P}^{*},$
    where $M_{2}\in\mathbb{R}^{mq\times{}mq}$.
@@ Appendix A Proof of 3
Claim ($i$) follows immediately from the implication
$Y\succeq{}0\implies{}A\left(Y\right)\succeq{}0.$
For ($ii$), we start by developing a formula for the dual cone. Let
$B\left(X\right)=A\left(TXT^{\top}\right).$
By Farkas’ lemma,
$\mathcal{P}^{*}=\left\{B^{*}\left(Y\right):Y\succeq{}0\right\}.$
@@
$\operatorname{tr}\left(B\left(X\right)Y\right)$ $=\operatorname{tr}\left(A\left(TXT^{\top}{}\right)A\left(Y\right)\right)$ (12)
@@
This shows that
$B^{*}\left(Y\right)=T^{\top}A\left(Y\right)T.$
It then follows that if
@@
$\begin{bmatrix}Y_{1}&Y_{2}\\ Y_{2}^{\top}&Y_{3}\end{bmatrix}=\begin{bmatrix}\operatorname{diag}_{m}\left(Y_{1}\right)&\operatorname{diag}_{m}\left(Y_{2}\right)\\ \operatorname{diag}\left(Y_{2}^{\top}\right)&\operatorname{diag}_{m}\left(Y_{3}\right)\end{bmatrix}\succeq{}0$
@@
$\begin{bmatrix}0&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}=T^{\top}\begin{bmatrix}0&Y_{2}\\ Y_{2}^{\top}&Y_{3}\end{bmatrix}T.$
@@
$\begin{bmatrix}M_{1}M_{2}^{+}M_{1}^{\top}&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}=T^{\top}\begin{bmatrix}Y_{2}Y_{3}^{+}Y_{2}^{\top}&Y_{2}\\ Y_{2}^{\top}&Y_{3}\end{bmatrix}T,$
which is in $\mathcal{P}^{*}$ as required. ∎
We now prove 3.
Proof.
Observe that by 1($i$), the optimization in 3 is lower bounded by the relaxed optimization problem
over $\{U_{1}(t),U_{2}(t)\}_{t=0}^{\infty},$
$\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\in\mathcal{P}_{Q},X(0)\succeq{}0.$
An analogous argument to that made in section 3.1 and 1($ii$) then shows that the set
$\mathcal{C}_{\mu}=\left\{\Lambda:\begin{bmatrix}\Lambda&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\in\mathcal{P}_{Q}^{*}\right\}$
is non-empty, it has a minimum element. Therefore under (a) and (b) in the theorem statement and by 2, the relaxed optimization problem meets all the conditions of 1, from which it soon follows that the relaxed optimization has a finite value if and only if there exists a
$\begin{bmatrix}\Lambda^{*}&0\\ 0&0\end{bmatrix}\in\mathcal{P}_{Q}^{*}$
such that
$\Lambda^{*}=S+F^{\top}\Lambda^{*}F-\left(F^{\top}\Lambda^{*}G+R_{1}\right)\left(R_{2}+G^{\top}{}\Lambda^{*}G\right)^{+}\left(G^{\top}\Lambda^{*}F+R_{1}^{\top}{}\right).$
Again by 1, the elements of the optimal input sequence $\overline{U}_{1}(t)$ and $\overline{U}_{2}(t)$ are the minimizers of the following optimization problem
Minimize $\operatorname{tr}\left(\begin{bmatrix}0&R_{1}+F^{\top}\Lambda^{*}G\\ R_{1}^{\top}+G^{\top}\Lambda^{*}F&R_{2}+G^{\top}\Lambda^{*}G\end{bmatrix}\begin{bmatrix}0&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\right)$
over $U_{1},U_{2},$
subject to $\begin{bmatrix}X(t)&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\in\mathcal{P}_{Q}.$
The result will then follow if
$\begin{bmatrix}X(t)&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\succeq{}0,$
since then all inputs to the relaxed problem are feasible also for the unrelaxed problem. It follows from an argument analogous to that in eq. 12, and from the assumption that $X(0)\succeq{}0$, that we may replace the constraint in the above with
$\begin{bmatrix}X(0)&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\succeq{}0$ (13)
without affecting the optimal value. Just as for the standard LQR problem it then follows from complementary slackness that there exists a $K$ such that
$\begin{bmatrix}I\\ K^{\top}\end{bmatrix}\Lambda^{*}\begin{bmatrix}I&K\end{bmatrix}\in\mathcal{P}_{Q}^{*}$
and
$\begin{bmatrix}I\\ K^{\top}\end{bmatrix}\Lambda^{*}\begin{bmatrix}I&K\end{bmatrix}\begin{bmatrix}X(0)&\overline{U}_{1}\\ \overline{U}_{1}^{\top}&\overline{U}_{2}\end{bmatrix}=0.$ (14)
Applying this control results in $X(1)\succeq{}0$, and rerunning this argument for all $t$ then shows that eqs. 13 and 14 hold for all $t$, not just $t=0$. Hence the optimal actions for the relaxed problem are feasible also for the unrelaxed problem, from which the result follows. ∎

### v2

@@ 3.2 A Linear Regulator on a Polyhedral Cone
In this subsection we will recover the results of [4, Theorem 1]. Consider the application of 1 when $\mathcal{P}$ is the polyhedral cone
@@
When minimizing the cost in 1($i$), we are in effect optimizing performance under the constraint that the control keeps the system state positive.
@@
It is readily checked that 1) amounts to requiring that $A\geq{}\left|{EB}\right|$. For 2) and 3) it is convenient to appeal to 2. From eq. 11 we see that for any $\mu$, $y$ and $z$ can be chosen according to
@@
where $w\geq{}0$. The set in 2($ii$) therefore has a minimum element given by $\overline{\lambda}=E^{\top}\left|{\mu}\right|$. This implies that for (2) we require that $s>E^{\top}\left|{r}\right|$, and also that
@@
3.3 A Structured Linear Quadratic Regulation Problem
@@
The following lemma establishes the key properties of the cone $\mathcal{P}_{Q}$ that we will require to apply 1. The first part of the result gives a formula for the dual cone, and establishes that a suitable function $\phi\left(\mu\right)$ exists (the function is in fact the same as that from section 3.1). The final part gives a formula that can be used to compute the optimal control law, and it is from this formula that we can deduce that the optimal control inherits additional structural features as claimed above.
Lemma 1.
Let $A$, $T$, $\mathcal{P}_{Q}$ and all the block partitioning of matrices be as in 1. Then
$\mathcal{P}_{Q}^{*}=\left\{T^{\top}A\left(Y\right)T:Y\succeq{}0\right\}.$
Furthermore if there exists a $\Lambda$ such that
$\begin{bmatrix}\Lambda&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\in\mathcal{P}^{*}_{Q},$ (12)
then for fixed $X$
$\inf\!\left\{\!\operatorname{tr}\!\left(\!\begin{bmatrix}0&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\!\!\begin{bmatrix}0&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\!\right)\!:\!\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\!\in\mathcal{P}_{Q}\!\right\}$ (13)
equals $\operatorname{tr}\left(M_{1}M_{2}^{+}M_{1}^{\top}X\right)$. A pair $\left(U_{1},U_{2}\right)$ that achieves the infimum above can be computed through the equation
$\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}=\begin{bmatrix}I\\ -M_{2}^{+}M_{1}^{\top}\end{bmatrix}X\begin{bmatrix}I\\ -M_{2}^{+}M_{1}^{\top}\end{bmatrix}^{\top}{}.$ (14)
@@
To understand how to leverage 1, consider again the optimal control problem from section 3.1, but suppose that the matrices $F$ and $G$ have a block-wise circulant structure. That is,
@@
It follows from these expressions that for such matrices,
$\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\in\mathcal{P}_{Q}\implies\begin{bmatrix}X(t+1)&0\\ 0&0\end{bmatrix}\in\mathcal{P}_{Q},$
where
$X(t+1)=\begin{bmatrix}F&G\end{bmatrix}\begin{bmatrix}X(t)&U_{1}(t)\\ U_{1}(t)^{\top}{}&U_{2}\left(t\right)\end{bmatrix}\begin{bmatrix}F^{\top}{}\\ G^{\top}\end{bmatrix}.$
Taking an analogous approach to that in section 3.1, we see that under suitable assumptions on the cost function, 1 can be applied to solve a relaxed version of the standard LQR problem, where the state and inputs can be drawn from $\mathcal{P}_{Q}$. Since $\mathcal{P}_{Q}$ contains the positive semi-definite cone, the solution to this problem will always give a lower bound on the achievable cost to the standard LQR problem. However more is true. In particular we see from the final part of 1 and eq. 14 in 1 that the optimal inputs to the relaxed problem can be computed through the control law $X\mapsto{}\left(U_{1},U_{2}\right)$ defined implicitly by
$\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}=\begin{bmatrix}I\\ -K\end{bmatrix}X\begin{bmatrix}I\\ -K\end{bmatrix}^{\top}{},$
where
$K=\left(R_{2}+G^{\top}{}\Lambda_{*}G\right)^{-1}\left(G^{\top}\Lambda_{*}F+R_{1}^{\top}{}\right).$
It follows from these expressions that whenever the initial condition satisfies $X\left(0\right)\succeq{}0$,
$\begin{bmatrix}X\left(t\right)&U_{1}\left(t\right)\\ U_{1}\left(t\right)^{\top}&U_{2}\left(t\right)\end{bmatrix}\succeq{}0$
for all $t$. Hence the found controls are also feasible, and hence also optimal, for the standard LQR problem. It is then easily seen from the expression for the dual cone that
$\left(Q^{-1}\!\otimes{}I_{q}\right)K\left(Q\otimes{}I_{p}\right)\!=\!\operatorname{diag}_{m}\left(\left(Q^{-1}\!\otimes{}I_{q}\right)K\left(Q\otimes{}I_{p}\right)\right).$
Therefore the optimal control law additionally inherits the block-wise circulant structure present in the problem data. Interestingly this cone based treatment makes no use of any orthogonality properties of the matrix $Q$ allowing similar insights to be derived for general dynamics that take the form of eq. 15, see for example [11].
@@ 4 Conclusions
We have proved a general result on optimal control on positive cones, covering standard linear quadratic control and corresponding results for positive systems as special cases. Of course, many other special cases can be derived by defining other relevant cones, or combinations of the ones above. Hopefully, this will in the future form the basis for a powerful toolbox combining the versatility of linear quadratic control with the rich scalability properties of positive systems and shortest path problems in networks. Preliminary steps in this direction have been taken in [12] and [13].
5 Acknowledgment
The authors are grateful to Yuchao Li for pointing out an error in the original statement of the main result.
Appendix A Proof of 1
@@
We start by developing the formula for the dual cone. Let $B\left(X\right)=A\left(TXT^{\top}\right)$. By Farkas’ lemma,
$\mathcal{P}^{*}_{Q}=\left\{B^{*}\left(Y\right):Y\succeq{}0\right\}.$
@@
$\operatorname{tr}\left(B\left(X\right)Y\right)$ $=\operatorname{tr}\left(A\left(TXT^{\top}{}\right)A\left(Y\right)\right)$
@@
This shows that $B^{*}\left(Y\right)=T^{\top}A\left(Y\right)T$ as required. To show the second claim observe that if
@@
$\begin{bmatrix}Y_{1}&Y_{2}\\ Y_{2}^{\top}&Y_{3}\end{bmatrix}=\begin{bmatrix}\operatorname{diag}_{m}\left(Y_{1}\right)&\operatorname{diag}_{m}\left(Y_{2}\right)\\ \operatorname{diag}_{m}\left(Y_{2}^{\top}\right)&\operatorname{diag}_{m}\left(Y_{3}\right)\end{bmatrix}\succeq{}0$
@@
$\begin{bmatrix}0&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}=T^{\top}\begin{bmatrix}0&Y_{1}\\ Y_{1}^{\top}&Y_{2}\end{bmatrix}T.$
@@
$\begin{bmatrix}M_{1}M_{2}^{+}M_{1}^{\top}&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}=T^{\top}\begin{bmatrix}Y_{2}Y_{2}^{+}Y_{1}^{\top}&Y_{1}\\ Y_{1}^{\top}&Y_{2}\end{bmatrix}T,$
which is in $\mathcal{P}^{*}_{Q}$. Since $\mathcal{P}_{Q}^{*}$ is contained in the semi-definite cone, $M_{1}M_{2}^{+}M_{1}^{\top}$ is the minimum element of the set
$\mathcal{C}_{\mu}=\left\{\Lambda:\begin{bmatrix}\Lambda&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\in\mathcal{P}^{*}_{Q}\right\},$
and so the formula for the infimum follows by 2. In order to verify the equation for the minimizer, first observe that from the above,
$A\left(T\begin{bmatrix}I\\ -M_{2}^{+}M_{1}^{\top}\end{bmatrix}X\begin{bmatrix}I\\ -M_{2}^{+}M_{1}^{\top}\end{bmatrix}^{\top}{}T^{\top}\right)=$
$\quad\begin{bmatrix}I&0\\ -Y_{2}^{+}Y_{1}^{\top}{}&0\end{bmatrix}A\left(T\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}T^{\top}\right)\begin{bmatrix}I&0\\ -Y_{2}^{+}Y_{1}^{\top}{}&0\end{bmatrix}^{\top}\!\!\!\!.$
It then follows that
$\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\in\mathcal{P}_{Q}\implies\begin{bmatrix}I\\ -M_{2}^{+}M_{1}^{\top}\end{bmatrix}X\begin{bmatrix}I\\ -M_{2}^{+}M_{1}^{\top}\end{bmatrix}^{\top}{}\!\!\!\!\!\!\in\mathcal{P}_{Q}.$
Next observe that since $\left(I-M_{2}^{+}M_{2}\right)M_{1}^{\top}=0$ (else there would exist no $\Lambda$ such that eq. 12 holds),
$\begin{bmatrix}M_{1}M_{2}^{+}M_{1}^{\top}\!\!&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\!\begin{bmatrix}I\\ -M_{2}^{+}M_{1}^{\top}\end{bmatrix}\!X\!\begin{bmatrix}I\\ -M_{2}^{+}M_{1}^{\top}\end{bmatrix}^{\top}{}\!\!\!\!\!\!=0.$ (16)
Finally for fixed $X$ note that the minimizer for eq. 13 equals the minimizer for
$\inf\!\left\{\!\operatorname{tr}\!\left(\!\begin{bmatrix}M_{1}M_{2}^{+}M_{1}^{\top}\!\!\!&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\!\!\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\!\right)\!:\!\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\!\in\mathcal{P}_{Q}\!\right\}\!.$
From the definition of a dual cone,
$\operatorname{tr}\!\left(\!\begin{bmatrix}M_{1}M_{2}^{+}M_{1}^{\top}&M_{1}\\ M_{1}^{\top}&M_{2}\end{bmatrix}\begin{bmatrix}X&U_{1}\\ U_{1}^{\top}&U_{2}\end{bmatrix}\right)\geq{}0.$
Therefore by eq. 16 the right hand side of eq. 14 is a minimizer as required. ∎